OpenStudy (anonymous):

Please help! The sum of two numbers is 56. If the smaller number is 1/3 the larger number, what are the two numbers? If x is the smaller number, one equation is...?

OpenStudy (anonymous):

Okay, so x is the number, y is the other x = 1/3 y x+y = 56 y =3x now use substitution to solve the system x+y = 56 x + 3x = 56 4x = 56 x = 14 so the smaller number, x, is 14.

OpenStudy (anonymous):

the other number is 14*3 = 42

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

np (: glad I could help